Is -0.002 irrational and a integer?
Answer by jim_thompson5910(35256)

-0.002 = -2/1000 = -1/500, which makes it a rational number. So it is NOT irrational.


It's also NOT an integer since it has a decimal part (that is nonzero). In other words, it is NOT a whole number (since it's NOT an integer).